OpenStudy (hannahduhe):

which of the following lengths are the sides of a right triangle? A.2.4, 3.2, 4 B.4.3, 4.4, 4.5 C. 3.1, 4.2, 4.8 D.All of the above

OpenStudy (johnweldon1993):

They would need to satisfy the Pythagorean Theorem \[\large a^2 + b^2 = c^2\] where 'a' and 'b' are two legs of the right triangle "Shorter sides" and 'c' is the hypotenuse "Longest side"

Directrix (directrix):

Video explanation. How Do You Determine if a Triangle Is a Right Triangle if You Know its Sides? http://www.virtualnerd.com/algebra-1/radical-expressions-equations/right-triangle-determination-from-sides.php

OpenStudy (raffle_snaffle):

a^2 + b^2 = c^2

OpenStudy (hannahduhe):

@raffle_snaffle @Directrix @johnweldon1993 is A.) CORRECT

zepdrix (zepdrix):

yay good job

OpenStudy (hannahduhe):

thanks